# MoM: Linear Sequence Modeling with Mixture-of-Memories

### v3

@@ 1 Introduction
Attention mechanisms have made significant contributions to the field of artificial intelligence, advancing various modalities such as language, vision, audio, video, graphs, and even time series (Achiam et al., 2023; Team, 2023). The Transformer (Vaswani, 2017), known for its ability to capture long-range dependencies, has become a foundational architecture in this space. However, traditional Transformers encounter computational challenges due to their quadratic time complexity, $O(n^{2})$, with respect to sequence length $n$, making it difficult to scale to long sequences. To overcome this limitation, several linear sequence modeling methods have been proposed, including linear attention (Katharopoulos et al., 2020; Qin et al., 2023a; Li et al., 2025), state space modeling (Gu & Dao, 2024; Dao & Gu, 2024), and linear RNNs (Peng et al., 2024; Qin et al., 2024d), which offer $O(n)$ training complexity and $O(1)$ inference complexity. These approaches often reduce the input sequence to a fixed-size hidden space, collapsing the information into a single “memory state”. While these methods enhance efficiency, they face two main challenges: limited memory capacity and memory interference. When new information overwrites the single fixed-size memory state, previously stored representations may degrade, which negatively impacts its long-term memory performance on recall-intensive tasks.
We argue that the strong performance of Transformer models on recall-intensive tasks arises from their ability to avoid memory interference by maintaining independent key-value caches for each token, thus offering virtually unlimited memory capacity. In contrast, linear sequence modeling relies on extreme compression, consolidating all the input information into a single fixed-size memory state (Katharopoulos et al., 2020; Dao & Gu, 2024). This approach results in limited memory capacity and inherently leads to memory interference issues.
@@ 2 Preliminary
To reduce the time complexity of Transformer attention, various optimization techniques have been proposed. Linear Transformers (Katharopoulos et al., 2020) replace the softmax attention mechanism with dot-product of feature maps $\phi(\cdot)$:
@@
where $\bm{q}_{t},\bm{k}_{t},\bm{v}_{t}\in\mathbb{R}^{d}$. While the presence of the denominator may lead to numerical instability (Qin et al., 2024b) and the feature map can utilize an identity function, which we omit for simplicity. In perspective of memory, the formulation can also be written in a recurrent format:
@@
Building on the foundational concepts of linear attention and memory perspective, some recent advancements have focused on optimizing memory structure, including gated updates (Yang et al., 2023; Qin et al., 2024e; d) and memory capacity expansion (Peng et al., 2024; Qin et al., 2024d).
@@ 3.1 Motivation
Linear sequence models compress the entire sequence data into a fixed-size memory state. Despite numerous efforts to minimize information loss, such as introducing gating mechanisms and employing more precise control over memory modifications (Orvieto et al., 2023; De et al., 2024; Beck et al., 2024; Yang et al., 2023; Zhang et al., 2024), some degradation in this compression process is inevitable. Expanding the memory capacity has been shown to mitigate this issue to some extent, with studies indicating that increasing memory capacity can enhance model performance (Qin et al., 2024d; Peng et al., 2024).
@@ 3.2 MoM: Mixture-of-Memories
To address the challenge outlined above, we propose a novel approach for encoding multi-item memory such as theta-gamma oscillations (Lisman & Jensen, 2013), and concepts from Mixture-of-Experts (MoE) (Shazeer et al., 2017), where different experts handle specific tokens. In this approach, we leverage multiple memory states, each of which is selectively updated by different inputs. This increases the memory capacity and enables the model to retain diverse pieces of information by storing various types of inputs in separate memory states.
@@ 3.2.2 Linear Recurrent Memory Module
    The equation above represents the simplest form of memory update for clarity. Our approach is flexible and does not rely on a specific memory update mechanism. To enhance performance, we can incorporate mechanisms such as forget gates (Sun et al., 2023).
@@ 4.1 Experimental Setups
Models.  In our experiments, we employ the Gated DeltaNet (Yang et al., 2024) as the memory update mechanism in MoM. The model is configured with four memory states, two of which are activated at each time step, along with a shared memory.
Baselines.  We evaluate MoM against several linear recurrent models and Transformers, including RetNet (Sun et al., 2023), GLA (Yang et al., 2023), Gated DeltaNet (Yang et al., 2024), and Transformer++ (Touvron et al., 2023), which incorporates Rotary Position Embeddings (Su et al., 2024) and GLU (Shazeer, 2020) into the Transformer architecture. To ensure a fair comparison, we train all baseline models from scratch using the exact same number of tokens.
Training.  We follow the training procedure described by Yang et al. (2023), utilizing the SlimPajama dataset (Soboleva et al., 2023) sampled with 100B tokens and tokenized using the Mistral tokenizer (Jiang et al., 2023). We train models from scratch with parameter sizes of 380M and 1.3B, respectively. For the 380M models, we train on 15B tokens with a batch size of 0.5M tokens. More detailed training configuration is provided in Appendix C. We utilized publicly available pretrained weights from Zhang et al. (2024) with exactly same configuration  [^1] Models marked with an asterisk use open-source pretrained weights with identical training configurations..
@@ 4.2.1 Recall-intensive Tasks
To thoroughly assess our model’s proficiency in such scenarios, we test six recall-intensive tasks following Arora et al. (2024): FDA (Arora et al., 2023), SWDE (Arora et al., 2023; Lockard et al., 2019), SQuAD (Rajpurkar et al., 2018), NQ (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017) and Drop (Dua et al., 2019). These tasks are designed to challenge a model’s ability to perform context-based retrieval and comprehension.
@@ 4.2.2 Long Context Tasks
Assessing performance on long-context tasks is crucial for linear models, as it reflects their ability to handle long-range dependencies effectively. We evaluated our model’s comprehension of long contexts using the LongBench benchmark (Bai et al., 2024; Contributors, 2023). In Table 3, we present the average results across various categories, including summarization, few-shot learning, synthetic tasks, and code completion, along with the overall mean across all tasks. The complete detailed results are provided in Appendix I.
@@ 4.2.5 Length Extrapolation
We pretrained the models on the Slimpajama dataset with a 2K context length and conducted extrapolation experiments on various lengths using the Fineweb(Penedo et al., 2024) dataset. We extended the length to 32K to calculate perplexity (ppl). As shown in Fig 4, the Transformer model experienced a significant increase in ppl due to its poor extrapolation capability. Among the linear models, MoM achieved the best results.
@@ 4.2.6 Memory Analysis
Memory Load Balance Analysis. To evaluate whether each memory segment in MoM is effectively balanced during inference on downstream tasks, we analyzed the number of tokens routed to each layer using around 300k tokens from the ARC-easy benchmark. We visualized the results with auxiliary loss (following the formulation introduced in Switch Transformer Fedus et al. (2022)) in Fig 5 with heatmaps and we also visualized results with auxiliary loss in Fig 9. Due to the adoption of auxiliary loss, the memory segments in each layer are almost uniformly routed and activated.
Functions of Different Memories. We analyzed the memory routing within the model and identified the categories of tokens processed by each of the four memories. By examining an intermediate layer, we observed that, unlike traditional MoE in FFN layers where experts often lack clear specialization and focus on syntactic punctuation(Jiang et al., 2023), MoM displays specialization among its memories. This finding suggests potential for exploring an MoE architecture where each memory serves a specific role. These distinctions among different memories are presented in Table 5.
@@ Appendix A Related Work
Linear recurrent models, comprising linear attention, linear RNNs, state-space models (SSMs), have garnered significant research interests (Qin et al., 2023b). The advancement of SSMs began with the pioneering work on S4 (Gu et al., 2022), which was later optimized through a diagonalized version (Gupta et al., 2022). Despite their strong performance on the LRA benchmark, these models have faced challenges in language modeling mainly because they rely solely on data-independent processes. As research progressed, constant forgetting gates were introduced, helping to alleviate some interference issues by uniformly managing memory decay (Sun et al., 2023; Gu & Dao, 2024). The next breakthrough involved data-dependent forget gates. These allowed models to dynamically adjust memory updates based on the input data, significantly enhancing performance across various tasks (Qin et al., 2024c; d; Yang et al., 2023; Zhang et al., 2024; Yang et al., 2024; Qin et al., 2024a). Sequence parallelism techniques are well adapted on linear recurrent models (Sun et al., 2024a; 2025) for efficient long context training. There have also been recent advancements in scaling law and test-time regression optimization (Shen et al., 2024; Sun et al., 2024c; Behrouz et al., 2024).
@@
Mixture-of-Experts (MoE) is a technique designed to enhance the capacity of deep neural networks while maintaining computational efficiency (Fedus et al., 2022; Rajbhandari et al., 2020; Lepikhin et al., 2020; Tang et al., 2023; Zhu et al., 2024; Qu et al., 2024). MoE achieves this by activating a subset of parameters, known as “experts”, for each input, which reduces the computational costs. Shazeer first integrated MoE into LSTM layers (Shazeer et al., 2017). The Switch Transformer (Fedus et al., 2022) refined this approach by simplifying the gating mechanism to select only one expert per input. Gshard (Lepikhin et al., 2020) further advanced this by using a top-2 expert routing strategy to improve performance. Recent MoE models, such as Deepseek-MoE (Dai et al., 2024), introduce shared experts to capture and consolidate common knowledge across different contexts, while designing fine-grained experts to increase combinatorial flexibility.
@@ Appendix B Comparison between MoM and MoE
  • Structure: In conventional MoE, each expert is a separate neural network within the feedforward network (FFN) layer such as Qwen-MoE (Team, 2024) and Linear-MoE Sun et al. (2024a). In contrast, in MoM, each memory is an RNN state with unique key-value projection weights to generate different key-value pairs. MoE operates during the channel mixing phase, where each token is processed independently by selected experts. On the other hand, MoM functions during the token mixing phase, where each memory processes different segments of the sequence, preserving inter-token relationships.
@@ Appendix C Experiments Details
For the 380M models, we train on 15B tokens with a batch size of 0.5M tokens. The warmup tokens count is set to 0.25M. We set the hidden ratio of our model to 3 to keep the activated parameter count approximately the same. For the 1.3B models, we train on 100B tokens with a batch size of 2M tokens. The warmup tokens count is 1B. We employ AdamW optimizer (Loshchilov et al., 2017; Sun et al., 2024b) with learning rate of 3e-4 with cosine learning rate schedule (Zhou et al., 2020). The weight decay is set to 0.01 and gradient clipping is 1.0. Our experiments were conducted using 32 NVIDIA A800 GPUs. Training the 380M parameter model required approximately 10 hours, while the 1.3B parameter model took around 6 days.
@@ Appendix D Commonsense Reasoning Tasks
As shown in Table 7, we report the language modeling perplexity and zero-shot performance of commonsense reasoning tasks following (Zhang et al., 2024) which includes WikiText (Merity et al., 2016), LAMBADA (Paperno et al., 2016), ARC-easy, ARC-challenge (Clark et al., 2018), HellaSwag (Zellers et al., 2019), PiQA (Bisk et al., 2020) and WinoGrande (Sakaguchi et al., 2019). The evaluation results are based on the lm-evaluation-harness (Gao et al., 2024).

### v4

@@ 1 Introduction
Attention mechanisms have made significant contributions to the field of artificial intelligence, advancing various modalities such as language, vision, audio, video, graphs, and even time series (achiam2023gpt; team2023internlm). The Transformer (vaswani2017attention), known for its ability to capture long-range dependencies, has become a foundational architecture in this space. However, traditional Transformers encounter computational challenges due to their quadratic time complexity, $O(n^{2})$, with respect to sequence length $n$, making it difficult to scale to long sequences. To overcome this limitation, several linear sequence modeling methods have been proposed, including linear attention (katharopoulos2020transformers; qin2023transnormerllm; li2025minimax), state space modeling (gu2024mambalineartimesequencemodeling; dao2024transformers), and linear RNNs (peng2024eagle; qin2024hgrn2), which offer $O(n)$ training complexity and $O(1)$ inference complexity. These approaches often reduce the input sequence to a fixed-size hidden space, collapsing the information into a single “memory state”. While these methods enhance efficiency, they face two main challenges: limited memory capacity and memory interference. When new information overwrites the single fixed-size memory state, previously stored representations may degrade, which negatively impacts its long-term memory performance on recall-intensive tasks.
We argue that the strong performance of Transformer models on recall-intensive tasks arises from their ability to avoid memory interference by maintaining independent key-value caches for each token, thus offering virtually unlimited memory capacity. In contrast, linear sequence modeling relies on extreme compression, consolidating all the input information into a single fixed-size memory state (katharopoulos2020transformers; dao2024transformers). This approach results in limited memory capacity and inherently leads to memory interference issues.
@@ 2 Preliminary
To reduce the time complexity of Transformer attention, various optimization techniques have been proposed. Linear Transformers (katharopoulos2020transformers) replace the softmax attention mechanism with dot-product of feature maps $\phi(\cdot)$:
@@
where $\bm{q}_{t},\bm{k}_{t},\bm{v}_{t}\in\mathbb{R}^{d}$. While the presence of the denominator may lead to numerical instability (qin2024lightning) and the feature map can utilize an identity function, which we omit for simplicity. In perspective of memory, the formulation can also be written in a recurrent format:
@@
Building on the foundational concepts of linear attention and memory perspective, some recent advancements have focused on optimizing memory structure, including gated updates (yang2023gated; qin2024hierarchically; qin2024hgrn2) and memory capacity expansion (peng2024eagle; qin2024hgrn2).
@@ 3.1 Motivation
Linear sequence models compress the entire sequence data into a fixed-size memory state. Despite numerous efforts to minimize information loss, such as introducing gating mechanisms and employing more precise control over memory modifications (orvieto2023resurrecting; de2024griffin; beck2024xlstm; yang2023gated; zhang2024gated), some degradation in this compression process is inevitable. Expanding the memory capacity has been shown to mitigate this issue to some extent, with studies indicating that increasing memory capacity can enhance model performance (qin2024hgrn2; peng2024eagle).
@@ 3.2 MoM: Mixture-of-Memories
To address the challenge outlined above, we propose a novel approach for encoding multi-item memory such as theta-gamma oscillations (lisman2013theta), and concepts from Mixture-of-Experts (MoE) (shazeer2017outrageously), where different experts handle specific tokens. In this approach, we leverage multiple memory states, each of which is selectively updated by different inputs. This increases the memory capacity and enables the model to retain diverse pieces of information by storing various types of inputs in separate memory states.
@@ 3.2.2 Linear Recurrent Memory Module
    The equation above represents the simplest form of memory update for clarity. Our approach is flexible and does not rely on a specific memory update mechanism. To enhance performance, we can incorporate mechanisms such as forget gates (sun2023retentive).
@@ 4.1 Experimental Setups
Models.  In our experiments, we employ the Gated DeltaNet (yang2024gated) as the memory update mechanism in MoM. The model is configured with four memory states, two of which are activated at each time step, along with a shared memory.
Baselines.  We evaluate MoM against several linear recurrent models and Transformers, including RetNet (sun2023retentive), GLA (yang2023gated), Gated DeltaNet (yang2024gated), and Transformer++ (touvron2023llama), which incorporates Rotary Position Embeddings (su2024roformer) and GLU (shazeer2020glu) into the Transformer architecture. To ensure a fair comparison, we train all baseline models from scratch using the exact same number of tokens.
Training.  We follow the training procedure described by yang2023gated, utilizing the SlimPajama dataset (cerebras2023slimpajama) sampled with 100B tokens and tokenized using the Mistral tokenizer (jiang2023mistral). We train models from scratch with parameter sizes of 380M and 1.3B, respectively. For the 380M models, we train on 15B tokens with a batch size of 0.5M tokens. More detailed training configuration is provided in Appendix C. We utilized publicly available pretrained weights from zhang2024gated with exactly same configuration  [^1] Models marked with an asterisk use open-source pretrained weights with identical training configurations..
@@ 4.2.1 Recall-intensive Tasks
To thoroughly assess our model’s proficiency in such scenarios, we test six recall-intensive tasks following arora2024simple: FDA (arora2023language), SWDE (arora2023language; lockard-etal-2019-openceres), SQuAD (rajpurkar2018know), NQ (kwiatkowski2019natural), TriviaQA (joshi2017triviaqa) and Drop (dua2019drop). These tasks are designed to challenge a model’s ability to perform context-based retrieval and comprehension.
@@ 4.2.2 Long Context Tasks
Assessing performance on long-context tasks is crucial for linear models, as it reflects their ability to handle long-range dependencies effectively. We evaluated our model’s comprehension of long contexts using the LongBench benchmark (bai2024longbenchbilingualmultitaskbenchmark; 2023opencompass). In Table 3, we present the average results across various categories, including summarization, few-shot learning, synthetic tasks, and code completion, along with the overall mean across all tasks. The complete detailed results are provided in Appendix I.
@@ 4.2.5 Length Extrapolation
We pretrained the models on the Slimpajama dataset with a 2K context length and conducted extrapolation experiments on various lengths using the Fineweb(penedo2024the) dataset. We extended the length to 32K to calculate perplexity (ppl). As shown in Fig 4, the Transformer model experienced a significant increase in ppl due to its poor extrapolation capability. Among the linear models, MoM achieved the best results.
@@ 4.2.6 Memory Analysis
Memory Load Balance Analysis. To evaluate whether each memory segment in MoM is effectively balanced during inference on downstream tasks, we analyzed the number of tokens routed to each layer using around 300k tokens from the ARC-easy benchmark. We visualized the results with auxiliary loss (following the formulation introduced in Switch Transformer fedus2022switch) in Fig 5 with heatmaps and we also visualized results with auxiliary loss in Fig 9. Due to the adoption of auxiliary loss, the memory segments in each layer are almost uniformly routed and activated.
Functions of Different Memories. We analyzed the memory routing within the model and identified the categories of tokens processed by each of the four memories. By examining an intermediate layer, we observed that, unlike traditional MoE in FFN layers where experts often lack clear specialization and focus on syntactic punctuation(jiang2023mistral), MoM displays specialization among its memories. This finding suggests potential for exploring an MoE architecture where each memory serves a specific role. These distinctions among different memories are presented in Table 5.
@@ Appendix A Related Work
Linear recurrent models, comprising linear attention, linear RNNs, state-space models (SSMs), have garnered significant research interests (qin2023scaling). The advancement of SSMs began with the pioneering work on S4 (gu2022efficientlymodelinglongsequences), which was later optimized through a diagonalized version (NEURIPS2022_9156b0f6). Despite their strong performance on the LRA benchmark, these models have faced challenges in language modeling mainly because they rely solely on data-independent processes. As research progressed, constant forgetting gates were introduced, helping to alleviate some interference issues by uniformly managing memory decay (sun2023retentive; gu2024mambalineartimesequencemodeling). The next breakthrough involved data-dependent forget gates. These allowed models to dynamically adjust memory updates based on the input data, significantly enhancing performance across various tasks (qin2024various; qin2024hgrn2; yang2023gated; zhang2024gated; yang2024gated; qin2024unlocking). Sequence parallelism techniques are well adapted on linear recurrent models (sun2024linear; sun2025lasp) for efficient long context training. There have also been recent advancements in scaling law and test-time regression optimization (shen2024scaling; sun2024learning; behrouz2024titans).
@@
Mixture-of-Experts (MoE) is a technique designed to enhance the capacity of deep neural networks while maintaining computational efficiency (fedus2022switch; rajbhandari2020zero; lepikhin2020gshard; tang2023ms; zhu2024llama; qu2024llama). MoE achieves this by activating a subset of parameters, known as “experts”, for each input, which reduces the computational costs. Shazeer first integrated MoE into LSTM layers (shazeer2017outrageously). The Switch Transformer (fedus2022switch) refined this approach by simplifying the gating mechanism to select only one expert per input. Gshard (lepikhin2020gshard) further advanced this by using a top-2 expert routing strategy to improve performance. Recent MoE models, such as Deepseek-MoE (dai2024deepseekmoe), introduce shared experts to capture and consolidate common knowledge across different contexts, while designing fine-grained experts to increase combinatorial flexibility.
@@ Appendix B Comparison between MoM and MoE
  • Structure: In conventional MoE, each expert is a separate neural network within the feedforward network (FFN) layer such as Qwen-MoE (qwen_moe) and Linear-MoE sun2024linear. In contrast, in MoM, each memory is an RNN state with unique key-value projection weights to generate different key-value pairs. MoE operates during the channel mixing phase, where each token is processed independently by selected experts. On the other hand, MoM functions during the token mixing phase, where each memory processes different segments of the sequence, preserving inter-token relationships.
@@ Appendix C Experiments Details
For the 380M models, we train on 15B tokens with a batch size of 0.5M tokens. The warmup tokens count is set to 0.25M. We set the hidden ratio of our model to 3 to keep the activated parameter count approximately the same. For the 1.3B models, we train on 100B tokens with a batch size of 2M tokens. The warmup tokens count is 1B. We employ AdamW optimizer (loshchilov2017fixing; sun2024co2) with learning rate of 3e-4 with cosine learning rate schedule (zhou2020pbsgd). The weight decay is set to 0.01 and gradient clipping is 1.0. Our experiments were conducted using 32 NVIDIA A800 GPUs. Training the 380M parameter model required approximately 10 hours, while the 1.3B parameter model took around 6 days.
@@ Appendix D Commonsense Reasoning Tasks
As shown in Table 7, we report the language modeling perplexity and zero-shot performance of commonsense reasoning tasks following (zhang2024gated) which includes WikiText (merity2016pointer), LAMBADA (paperno2016lambada), ARC-easy, ARC-challenge (Clark2018ThinkYH), HellaSwag (zellers2019hellaswag), PiQA (Bisk2020) and WinoGrande (sakaguchi2019winogrande). The evaluation results are based on the lm-evaluation-harness (eval-harness).